# SlowFast-LLaVA: A strong training-free baseline for video large language models

### v1

@@ Abstract
We propose SlowFast-LLaVA (or SF-LLaVA for short), a training-free video large language model (LLM) that can jointly capture the detailed spatial semantics and long-range temporal context without exceeding the token budget of commonly used LLMs. This is realized by using a two-stream SlowFast design of inputs for Video LLMs to aggregate features from sampled video frames in an effective way. Specifically, the Slow pathway extracts features at a low frame rate while keeping as many spatial details as possible (e.g., with $24\times 24$ tokens), and the Fast pathway operates on a high frame rate but uses a larger spatial pooling stride (e.g., downsampling 6×6$\times$6 ×) to focus on the motion cues. As a result, this design allows us to adequately capture both spatial and temporal features that are beneficial for understanding details along the video. Experimental results show that SF-LLaVA outperforms existing training-free methods on a wide range of video tasks. On some benchmarks, it achieves comparable or even better performance compared to state-of-the-art Video LLMs that are fine-tuned on video datasets.
Figure 1: Comparison with state-of-the-art 7B Video LLMs on 8 video benchmarks. Training-free and supervised fine-tuned (SFT) Video LLMs are marked using solid (—) and dashed (- - -) lines, respectively. SF-LLaVA outperforms existing training-free methods on all benchmarks, and achieves even superior results compared to most SFT methods fine-tuned on video data.
@@ 1 Introduction
Video large language models (LLMs) process video inputs and generate coherent and contextually relevant responses to user commands by applying a pre-trained LLM (Achiam et al., 2023; Chiang et al., 2023; Touvron et al., 2023b; Jiang et al., 2024). Although achieving convincing results, Video LLMs (Maaz et al., 2024b; Lin et al., 2023; Xu et al., 2024; Zhang et al., 2024b) typically are trained on supervised fine-tuning (SFT) video datasets, leading to high computational and labeling cost. Recently, training-free methods (Kim et al., 2024; Wu, 2024; Zhang et al., 2024b) have been proposed as a simple and highly cost-efficient solution. They directly use well-trained Image LLMs for video tasks without any additional fine-tuning and demonstrate convincing performance. However, most existing Video LLMs have two main drawbacks: (1) they work effectively only with a limited number of frames as inputs (e.g., 6 for IG-VLM (Kim et al., 2024) and 16 for PLLaVA (Xu et al., 2024)), making it difficult to capture fine-grained spatial and temporal content throughout the video, and (2) they simply feed the video features into an LLM without a proper temporal modeling design and fully rely on the capability of the LLM to model the motion patterns.
@@
We present SlowFast-LLaVA (or SF-LLaVA for short), a training-free Video LLM that is built upon LLaVA-NeXT (Liu et al., 2024) without further fine-tuning. Inspired by the successful two-stream networks (Simonyan & Zisserman, 2014; Feichtenhofer et al., 2019) for action recognition, we propose a new SlowFast design of inputs for Video LLMs to capture both detailed spatial semantics and long-range temporal context. Specifically, the Slow pathway extracts features at a low frame rate while keeping the spatial details at a higher resolution (e.g., 8 frames each with $24\times 24$ tokens), and the Fast pathway operates on a high frame rate but uses an aggressive spatial pooling stride (e.g., downsampling each frame to $4\times 4$ tokens) to focus on motion cues. SF-LLaVA combines the “Slow and Fast features” together as an effective representation for various video tasks. SF-LLaVA has two main advantages over prior work. First, it integrates complementary features from the slowly changing visual semantics and rapidly changing motion dynamics, providing a comprehensive understanding of videos. Second, the dual-pathway design balances the modeling capability and computational efficiency, and enables us to input more video frames to preserve adequate details.
@@
SF-LLaVA takes a video as input by uniformly sampling a large number of video frames ($N$) to maintain its detailed spatial and temporal information. Frame features $\mathbf{F}_{v}$ are extracted independently via a visual encoder (e.g., CLIP-L (Radford et al., 2021)) followed by a visual-language adaptor for feature alignment. Then, the features $\mathbf{F}_{v}$ are fed into the Slow and Fast pathways separately. The Slow pathway will uniformly sample $N^{\textnormal{slow}}\ll N$ features from $\mathbf{F}_{v}$. Prior work (Xu et al., 2024) found that properly pooling frame features can improve both the efficiency and robustness. We follow them to aggregate features in the Slow pathway by using a pooling with a small stride (e.g., $1\times 2$) over the spatial dimensions. The Fast pathway takes all $N^{\textnormal{fast}}=N$ features and performs a more aggressive spatial pooling of each frame to focus on a finer temporal resolution. Visual tokens from both pathways are then concatenated and fed into the LLM to get an answer.
@@
We extensively evaluate SF-LLaVA on 3 different video tasks (i.e., Open-Ended VideoQA, Multiple Choice VideoQA, and Text Generation) with 8 benchmarks including videos from various scenarios (e.g., first- and third-person views) and types (e.g., short and long videos). Experimental results (as shown in Fig. 1) show that SF-LLaVA outperforms existing training-free methods by a clear margin on all benchmarks, and achieves on-par or even better performance compared to state-of-the-art SFT models that have been carefully fine-tuned on video datasets. We also conduct comprehensive ablation studies on our SlowFast design recipe, which hopefully provide some valuable insights for future work on Video LLMs.
@@ 2.1 Image Large Language Models
Significant advance has been observed in the computer vision area with the development of image large language models (LLMs) (Achiam et al., 2023; Team et al., 2023; McKinzie et al., 2024; Abdin et al., 2024; Liu et al., 2024). As a pioneer work, Flamingo (Alayrac et al., 2022) accepts arbitrarily interleaved visual and text data as inputs and generates text in an open-ended manner. BLIP-2 (Li et al., 2023b) uses pretrained visual and text models, and bridges the domain gap with the proposed Q-Former. LLaVA(-v1.5/NeXT) (Liu et al., 2023b; a; 2024) achieves remarkable performance by leveraging a simple linear connector or an MLP between visual and text models and designing an efficient instruction following data pipeline assisted with GPT. More recently, MM1 (McKinzie et al., 2024) conducts comprehensive ablation studies on model components and data choices, and offers valuable insights for understanding Image LLMs. There are also efforts on ingesting other modalities. Ferret (You et al., 2023; Zhang et al., 2024a) focuses on the box/shape modality and enhances a model’s language grounding capability at any granularity. 3D-LLM (Hong et al., 2023) enables open-ended question answering in 3D by injecting 3D representations into an LLM. 4M (Mizrahi et al., 2023; Bachmann et al., 2024) presents a general any (modality) to any (modality) framework with strong out-of-box perceptional and generative capabilities.
2.2 Video Large Language Models
@@
With the rapid development of LLMs (Achiam et al., 2023; Team et al., 2023; Chiang et al., 2023; Touvron et al., 2023a; b), there is increasing interest in generalist video models that can perform a wide range of video tasks. Video-ChatGPT (Maaz et al., 2024b) extracts per-frame features then aggregates them by using two spatial and temporal pooling operations before inputting them to an LLM. VideoChat (Li et al., 2023c) encodes a video as both video text descriptions and video appearance embeddings. Video-LLaVA (Lin et al., 2023) pre-aligns the image and video encoders, and learns a shared projector to project them to the language space. PLLaVA (Xu et al., 2024) achieves convincing performance by fine-tuning a pre-trained Image LLM on video understanding data. LLaVA-NeXT-Video (Zhang et al., 2024b) improves LLaVA-NeXT (Liu et al., 2024) by fine-tuning it on video data, and its DPO version (Zhang et al., 2024b) further aligns the model responses with AI feedbacks. Although achieving promising performance, all above methods need to be fine-tuned on large-scale video data, leading to high computation and labeling cost.
@@
Training-Free Video LLMs are built upon Image LLMs and require no additional fine-tuning to work for video scenarios. FreeVA (Wu, 2024) carefully explores different temporal aggregation strategies and effectively pools video frame features before sending them to an LLM. IG-VLM (Kim et al., 2024) assembles multiple video frames into an image grid and uses the Image LLM as it is on the image grid for answering video-related questions. These training-free models show surprisingly good results on various video benchmarks, but they have at least two main drawbacks. First, they simply ingest visual frame features and fully rely on the capability of the LLMs to capture the temporal dependency along the entire video. Second, they can only successfully process a few frames from a video (e.g., 4 frames in FreeVA and 6 frames in IG-VLM), which limits them to work only for short and simple videos. In this paper, we propose a new SlowFast design that uses two-stream inputs for Video LLMs. By effectively and efficiently consuming a large number of frames, this design allows us to jointly capture detailed spatial and temporal cues for video understanding.
@@ 3 SlowFast-LLaVA
We introduce a training-free Video LLM, named SlowFast-LLaVA (or SF-LLaVA for short), based on the LLaVA-NeXT (Liu et al., 2024) framework, as shown in Fig. 2. Inspired by (Simonyan & Zisserman, 2014; Feichtenhofer et al., 2019) for action recognition, we propose a SlowFast design that uses two-stream inputs for Video LLMs to jointly capture the detailed spatial semantic and long-range temporal context without exceeding the token budget of commonly used LLMs. (e.g., 4096 in Vicuna-v1.5). Specifically, the Slow pathway includes “high-resolution” [^1] We mean “high- or low-resolution” frames by their number of tokens after the visual encoder and pooling, such as $24\times 24$ or $4\times 4$, not the raw image size. We extract features for all frames in size of $336\times 336$. but low-frame-rate frame features (e.g., 8 frames each with $24\times 24$ tokens) to capture the detailed spatial semantics, and the Fast pathway includes “low-resolution” but high-frame-rate frame features (e.g., 64 frames each with $4\times 4$ tokens) to model greater temporal context. This design allows us to adequately preserve both spatial and temporal information, and aggregate them together as a powerful representation for video understanding. In this section, we will first introduce the background and pipeline of existing training-free Video LLMs, and then explain the details of SF-LLaVA’s input and architecture.
@@ 3.1 Preliminaries: Training-free Video LLMs
A training-free Video LLM is built upon a pre-trained Image LLM without further fine-tuning on any data. It saves significant computation resources and model training time, and offers a greater flexibility that can be quickly adapted into different image and video scenarios. The main effort of this research direction focuses on improving the video representation (e.g., organizing sampled frames (Kim et al., 2024) and incorporating textual descriptions (Zhang et al., 2023a)) and effectively leveraging the knowledge of a pre-trained LLM to better fit into the video tasks.
@@
Figure 2: Visualization of SlowFast-LLaVA, which is a training-free model built upon LLaVA-NeXT without further fine-tuning. The Slow pathway (in color yellow) extracts features at a low frame rate while keeping as many spatial details as possible with more tokens, and the Fast pathway (in color blue) operates on a high frame rate but applies a larger spatial pooling stride to focus on the motion cue. This design allows us to adequately preserve adequate spatial and temporal information, and aggregate them together as an effective representation for understanding details in videos.
@@ 3.2 SlowFast Architecture
As shown in Fig. 2, our SF-LLaVA follows the standard training-free Video LLM pipeline. It takes a video $\mathbf{V}$ and a question $\mathbf{Q}$ as inputs, and outputs an answer $\mathbf{A}$, in respond to $\mathbf{Q}$. For the input, we uniformly sample $N$ frames, $\mathbf{I}=\{I_{1},I_{2},...,I_{N}\}$, from each video in an arbitrary size and length, without special frame assembling. The video features are extracted frame by frame independently as $\mathbf{F}_{v}\in\mathbb{R}^{N\times H\times W}$, where $H$ and $W$ are the height and width of the frame feature. Then, we further process $\mathbf{F}_{v}$ in two streams (i.e., the Slow and Fast pathways as follows), and combine them together as an effective video representation.
@@
Finally, the aggregated video feature is obtained by $\mathbf{F}_{v}^{\textnormal{aggr}}=[\texttt{flat}(\mathbf{F}_{v}^{\textnormal{slow}}),\texttt{flat}(\mathbf{F}_{v}^{\textnormal{fast}})]$, where flat and [,][,][ , ] indicate the flatten and concatenation operations, respectively. As the equation implies, we do not use any special tokens in $\mathbf{F}_{v}^{\textnormal{aggr}}$ to separate the Slow and Fast pathway. Thus, SF-LLaVA uses $N^{\textnormal{slow}}\times H^{\textnormal{slow}}\times W^{\textnormal{slow}}+N^{\textnormal{fast}}\times H^{\textnormal{fast}}\times W^{\textnormal{fast}}$ video tokens in total.
@@ 4.1 Benchmarks and Metrics
We evaluate SF-LLaVA against state-of-the-art Video LLMs on three categories of zero-shot video question-answering (VideoQA) tasks, each consists of multiple datasets.
Open-Ended VideoQA expects the model to generate answers in a free style in respond to a question for a video. We include MSVD-QA (Chen & Dolan, 2011), MSRVTT-QA (Xu et al., 2016), TGIF-QA (Li et al., 2016) and ActivityNet-QA (or ANet-QA in tables) (Yu et al., 2019) as the benchmarks for this task. Except for ActivityNet-QA, we follow prior work (Maaz et al., 2024b) and report the performance on the validation set. We use the GPT-assisted evaluation to assess the accuracy (accuracy with answer true or false) and the quality (score ranging from 0 to 5) of the models. As pointed out by FreeVA (Wu, 2024) that different GPT versions can significantly impact the results, we report to use GPT-3.5-Turbo-0125 to perform a fair comparison.
@@
Text Generation is used to evaluate the text generation performance of a Video LLM, and especially focuses on the follow aspects: Correctness of Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (CO). We use the VCGBench (Maaz et al., 2024b) to evaluate these tasks and follow its official pipeline to evaluate this capability. Specifically, we use GPT-3.5-Turbo-0125 for evaluation.

### v2

@@ Abstract
We propose SlowFast-LLaVA (or SF-LLaVA for short), a training-free video large language model (LLM) that can jointly capture detailed spatial semantics and long-range temporal context without exceeding the token budget of commonly used LLMs. This is realized by using a two-stream SlowFast design of inputs for Video LLMs to aggregate features from sampled frames in an effective way. Specifically, the Slow pathway extracts features at a low frame rate while keeping as much spatial detail as possible (e.g., with $12\times 24$ tokens), and the Fast pathway operates on a high frame rate but uses a larger spatial pooling stride (e.g., downsampling 6×6$\times$6 ×) to focus on the motion cues. As a result, this design allows us to adequately capture both spatial and temporal features that are beneficial for detailed video understanding. Experimental results show that SF-LLaVA outperforms existing training-free methods on a wide range of video tasks. On some benchmarks, it achieves comparable or even better performance compared to state-of-the-art Video LLMs that are fine-tuned on video datasets. Code has been made available at: https://github.com/apple/ml-slowfast-llava.
Figure 1: Comparison with state-of-the-art 7B Video LLMs on 8 video benchmarks. Training-free and supervised fine-tuned (SFT) Video LLMs are marked using solid (—) and dashed (- - -) lines, respectively. SF-LLaVA outperforms existing training-free methods on all benchmarks, and achieves even better results compared to most SFT methods that are fine-tuned on video datasets.
@@ 1 Introduction
Video large language models (LLMs) process video inputs and generate coherent and contextually appropriate responses to user commands by using a pre-trained LLM (Achiam et al., 2023; Chiang et al., 2023; Touvron et al., 2023b; Jiang et al., 2024). Although achieving convincing results, most Video LLMs (Maaz et al., 2024b; Lin et al., 2023; Xu et al., 2024; Zhang et al., 2024b) are fine-tuned on large-scale labeled video datasets, leading to high computational and labeling cost. Recently, training-free methods (Kim et al., 2024; Wu, 2024; Zhang et al., 2024b) have been proposed as a simple and highly cost-efficient solution. They directly use well-trained Image LLMs for video tasks without additional fine-tuning and demonstrate encouraging performance. However, most existing Video LLMs have two main drawbacks: (1) they work effectively only with a limited number of frames as inputs (e.g., 6 for IG-VLM (Kim et al., 2024) and 16 for PLLaVA (Xu et al., 2024)), making them difficult to capture fine-grained spatial and temporal content throughout the video, and (2) they simply feed the video features into an LLM without a proper temporal modeling design and fully rely on the capability of the LLM to model the motion patterns.
@@
We present SlowFast-LLaVA (or SF-LLaVA for short), a training-free Video LLM that is built upon LLaVA-NeXT (Liu et al., 2024) without further fine-tuning. Inspired by the successful two-stream networks (Simonyan & Zisserman, 2014; Feichtenhofer et al., 2019) for action recognition, we propose a new SlowFast design of inputs for Video LLMs to capture both detailed spatial semantics and long-range temporal context. Specifically, the Slow pathway extracts features at a low frame rate while keeping spatial information at a higher resolution (e.g., 8 frames each with $24\times 24$ tokens), and the Fast pathway operates on a high frame rate but uses an aggressive spatial pooling stride (e.g., downsampling each frame to $4\times 4$ tokens) to focus on motion cues. SF-LLaVA combines the “Slow and Fast features” together as an effective video representation for various tasks. SF-LLaVA has two main advantages over prior work. First, it integrates complementary features from the slowly changing visual semantics and rapidly changing motion dynamics, providing a comprehensive understanding of videos. Second, the dual-pathway design balances the modeling capability and computational efficiency, and enables us to input more video frames to preserve adequate details.
@@
SF-LLaVA takes a video as input by uniformly sampling a large number of frames (denoted as $N$) to maintain as much detail as possible. Frame features $\mathbf{F}_{v}$ are extracted independently via a visual encoder (e.g., CLIP-L (Radford et al., 2021)) followed by a visual-language adaptor for feature alignment. Then, the features $\mathbf{F}_{v}$ are fed into the Slow and Fast pathways separately. The Slow pathway uniformly samples $N^{\textnormal{slow}}\ll N$ features from $\mathbf{F}_{v}$. Prior work (Xu et al., 2024) found that properly pooling frame features can improve both efficiency and robustness. We follow them to aggregate features in the Slow pathway by using a pooling with a small stride (e.g., $1\times 2$) over the spatial dimensions. The Fast pathway takes all $N^{\textnormal{fast}}=N$ features and performs a more aggressive spatial pooling of each frame to focus on a finer temporal resolution. Finally, visual tokens from both pathways are concatenated and fed into the LLM to generate the answer.
@@
We extensively evaluate SF-LLaVA on 3 video tasks (i.e., Open-Ended VideoQA, Multiple Choice VideoQA, and Text Generation) with 8 benchmark, including videos from various types (e.g., first- and third-person views) and lengths (e.g., short and long videos). Experimental results (as shown in Fig. 1) show that SF-LLaVA outperforms existing training-free methods by a clear margin on all benchmarks, and achieves on-par or even better performance compared to SFT models that have been carefully fine-tuned on video datasets. We also conduct comprehensive ablation studies on our SlowFast design recipe, which hopefully provide some valuable insights for future work.
@@ 2 Related Work
Image Large Language Models. Significant advances have been observed in the development of multimodal large language models (LLMs) (Achiam et al., 2023; Team et al., 2023; McKinzie et al., 2024; Abdin et al., 2024; Liu et al., 2024). As a pioneer work, Flamingo (Alayrac et al., 2022) accepts arbitrarily interleaved visual and text data as inputs and generates text in an open-ended manner. BLIP-2 (Li et al., 2023b) uses pre-trained visual and text models, and bridges the domain gap with the proposed Q-Former. LLaVA(-v1.5/NeXT) (Liu et al., 2023b; a; 2024) achieves remarkable performance by leveraging a simple linear connector or an MLP between visual and text models and designing an efficient instruction following data pipeline assisted with GPT. More recently, MM1 (McKinzie et al., 2024) conducts comprehensive ablation studies on model components and data choices, and offers valuable insights for understanding Image LLMs. There are also efforts to ingest other modalities. Ferret (You et al., 2023; Zhang et al., 2024a) focuses on the box/shape modality and enhances a model’s language grounding capability at any granularity. 3D-LLM (Hong et al., 2023) enables open-ended question answering in 3D by injecting 3D representations into an LLM. 4M (Mizrahi et al., 2023; Bachmann et al., 2024) presents a general any (modality) to any (modality) framework with strong out-of-box perceptional and generative capabilities.
@@
Video Large Language Models. With the rapid development of LLMs (Achiam et al., 2023; Team et al., 2023; Chiang et al., 2023; Touvron et al., 2023a; b), there is increasing interest in generalist video models that can perform a wide range of video tasks. Video-ChatGPT (Maaz et al., 2024b) extracts per-frame features then aggregates them by using two spatial and temporal pooling operations before inputting them to an LLM. VideoChat (Li et al., 2023c) encodes a video as both video text descriptions and video appearance embeddings. Video-LLaVA (Lin et al., 2023) pre-aligns the image and video encoders, and learns a shared projector to project them to the language space. PLLaVA (Xu et al., 2024) achieves convincing performance by fine-tuning a pre-trained Image LLM on video understanding data. LLaVA-NeXT-Video (Zhang et al., 2024b) improves LLaVA-NeXT (Liu et al., 2024) by fine-tuning it on video data, and its DPO version (Zhang et al., 2024b) further aligns the model responses with AI feedbacks.
@@
Training-Free Video LLMs are built upon Image LLMs and require no additional fine-tuning to work for video scenarios. FreeVA (Wu, 2024) explores different temporal aggregation methods and effectively pools video features before sending them to an LLM. IG-VLM (Kim et al., 2024) assembles multiple video frames into an image grid and uses the Image LLM as it is on the image grid for video tasks. These training-free models show encouraging results on various benchmarks, but they have two main drawbacks. First, they can only successfully process a few frames from a video (e.g., 4 frames in FreeVA and 6 frames in IG-VLM), which limits them to work only for short and simple videos. Second, they simply ingest the video features and fully rely on the capability of the LLMs to capture the temporal dependency along the video. In this paper, we propose a new SlowFast design to capture both detailed spatial and temporal cues for video understanding by effectively and efficiently taking more frames (e.g., 50) as inputs.
@@ 3 SlowFast-LLaVA
We introduce a training-free Video LLM, named SlowFast-LLaVA (or SF-LLaVA for short), based on the LLaVA-NeXT (Liu et al., 2024), as shown in Fig. 2. Inspired by (Simonyan & Zisserman, 2014; Feichtenhofer et al., 2019) for action recognition, we propose a SlowFast design that uses two-stream inputs for Video LLMs to jointly capture detailed spatial semantic and long-range temporal context without exceeding the token budget of commonly used LLMs. (e.g., 4096 in Vicuna-v1.5). Specifically, the Slow pathway includes “high-resolution” [^1] We mean “high- or low-resolution” frames by their number of tokens after the visual encoder and pooling, such as $24\times 24$ or $4\times 4$, not the raw image size. We extract features for all frames in size of $336\times 336$. but low-frame-rate frame features (e.g., 8 frames each with $12\times 24$ tokens) to capture spatial detail as much as possible, and the Fast pathway includes “low-resolution” but high-frame-rate frame features (e.g., 64 frames each with $4\times 4$ tokens) to model greater temporal context. This design allows us to adequately preserve both spatial and temporal information, and aggregate them together as a powerful video representation.
@@ 3.1 Preliminaries: Training-free Video LLMs
A training-free Video LLM is built upon a pre-trained Image LLM without further fine-tuning on any data. It saves significant computation resources and model training time, and offers a greater flexibility that can be quickly adapted into different application scenarios. The main effort of this research direction is to improve the visual representation (e.g., organizing sampled frames (Kim et al., 2024) or incorporating textual descriptions (Zhang et al., 2023a)) and effectively leveraging the knowledge of a pre-trained LLM to better fit into the video tasks.
@@
Figure 2: Visualization of SlowFast-LLaVA, which is a training-free model built upon LLaVA-NeXT without further fine-tuning. The Slow pathway (in color yellow) extracts features at a low frame rate while keeping as much spatial detail as possible with more tokens, and the Fast pathway (in color blue) operates on a high frame rate but applies a larger spatial pooling stride to focus on the motion cue. This design allows us to adequately preserve adequate spatial and temporal information, and aggregate them together as an effective representation for detailed video understanding.
@@ 3.2 SlowFast Architecture
As shown in Fig. 2, our SF-LLaVA follows the standard training-free Video LLM pipeline. It takes a video $\mathbf{V}$ and a question $\mathbf{Q}$ as inputs, and outputs an answer $\mathbf{A}$, in response to $\mathbf{Q}$. For the input, we uniformly sample $N$ frames, $\mathbf{I}=\{I_{1},I_{2},...,I_{N}\}$, from each video in an arbitrary size and length, without special frame assembling. The video features are extracted frame by frame independently as $\mathbf{F}_{v}\in\mathbb{R}^{N\times H\times W}$, where $H$ and $W$ are the height and width of the frame feature. Then, we further process $\mathbf{F}_{v}$ in two streams (i.e., the Slow and Fast pathways as follows), and combine them together as an effective video representation.
@@
Finally, the aggregated video feature is obtained by $\mathbf{F}_{v}^{\textnormal{aggr}}=[\texttt{flat}(\mathbf{F}_{v}^{\textnormal{slow}}),\texttt{flat}(\mathbf{F}_{v}^{\textnormal{fast}})]$, where flat and [,][,][ , ] indicate the flatten and concatenation operations, respectively. As the equation implies, we do not use any special tokens in $\mathbf{F}_{v}^{\textnormal{aggr}}$ to separate the Slow and Fast pathways. Thus, SF-LLaVA uses $N^{\textnormal{slow}}\times H^{\textnormal{slow}}\times W^{\textnormal{slow}}+N^{\textnormal{fast}}\times H^{\textnormal{fast}}\times W^{\textnormal{fast}}$ video tokens in total.
@@ 4.1 Benchmarks and Metrics
Open-Ended VideoQA expects the model to generate answers in freestyle in response to a question for a video. We include MSVD-QA (Chen & Dolan, 2011), MSRVTT-QA (Xu et al., 2016), TGIF-QA (Li et al., 2016) and ActivityNet-QA (or ANet-QA in tables) (Yu et al., 2019) as the benchmarks for this task. Except for ActivityNet-QA, we follow prior work (Maaz et al., 2024b) and report the performance on the validation set. We use the GPT-assisted evaluation to assess the accuracy (accuracy with the answer being true or false) and the quality (score ranging from 0 to 5) of the models. As pointed out by FreeVA (Wu, 2024) different GPT versions can significantly impact the results, we report to use GPT-3.5-Turbo-0125 to perform a fair comparison.
@@
Text Generation is used to evaluate the text generation performance of a Video LLM, and especially focuses on the following aspects: Correctness of Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (CO). We use the VCGBench (Maaz et al., 2024b) to evaluate these tasks and follow its official pipeline to evaluate this capability. Specifically, we use GPT-3.5-Turbo-0125 for evaluation.